Given that z= xysin(2x+3y) find (a) z_xx (b) z_yy


1
Expert's answer
2020-04-02T13:24:17-0400

(a) "z=xysin(2x+3y)"

differentiating with respect to x considering y as constant (using product rule)

"z_x=xy\cdot cos(2x+3y)\cdot2+sin(2x+3y)\cdot y"

"z_x=2xycos(2x+3y)+ysin(2x+3y)"

differentiating with respect to x once again

"z_{xx}=2xy\cdot (-sin(2x+3y))\cdot2+cos(2x+3y)\cdot2y+y\cdot cos(2x+3y)\cdot2"

"z_{xx}=-4xysin(2x+3y)+4ycos(2x+3y)"


(b) "z=xysin(2x+2y)"

differentiating with respect to y considering x as constant ( using product rule)

"z_y=xy\cdot cos(2x+3y)\cdot3+sin(2x+3y)\cdot x"

"z_y=3xycos(2x+3y)+xsin(2x+3y)"

differentiating with respect to y again

"z_{yy}=3xy \cdot(-sin(2x+3y)\cdot 3)+cos(2x+3y)\cdot3x+x\cdot cos(2x+3y)\cdot 3"

"z_{yy}=-9xysin(2x+3y)+6xcos(2x+3y)"